Consider the reaction  PCl3(g) + Cl2(g)  ⇋   PCl5(g)   Kc = 0.040 at 450oC. If a reaction is initiated with 0.40 mole of Cl2 and 0.40 mole of PCl3 in a 2.0 liter container, what is the equilibrium concentration of PCl5 in the same system?




1
Expert's answer
2021-12-07T06:31:01-0500

"K \n\nc\n\n\u200b\t\n\n =20= \n\n\\frac{(0.4+x)}{(0.2\u2212x)(0.3\u2212x)}"

"20= 

\frac{0.4+x}{0.06−0.5x+x 

^2}"


∴20x2 −11x+0.8=0


∴x=0.08